How much time does it take for the particle to accelerate from 68.2 m/s to 109.4 m/s?
1
Expert's answer
2012-10-09T08:37:38-0400
The time needed for the particle to accelerate from 68.2 m/s to 109.4 m/s depends on the acceleration of the particle, or which is the same, on the force acting on it and its mass. For example, if the mass of the particle is M and the force acting on it is F, then the acceleration of a particle is

a = F/M

and it takes

T = (109.4[m/s]-68.2[m/s])/a = 41.2/a seconds

to accelerate from 68.2 m/s to 109.4 m/s.